# GenKI: Enhancing Open-Domain Question Answering with Knowledge Integration and Controllable Generation in Large Language Models

### v1

@@ Abstract.
Cross-Domain Sequential Recommendation, Large Language Model
copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; Appril 28–May 02, 2025; Sydney, SYDisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Open-domain question answering (OpenQA) has garnered significant attention in the Natural Language Processing (NLP) community, as it offers enhanced user-friendliness recommendation (wang2025generative, ; guo2024scaling, ; wu2024survey, ; liu2023user, ) and efficiency compared to conventional search engines. OpenQA seeks to respond to questions utilizing auxiliary knowledge sources (zhu2021retrieving, ), necessitating models that possess both knowledge capability and comprehension abilities. Conventional methods usually entail a retriever-reader framework (DBLP:conf/iclr/DasDZM19, ). The evolution of retriever models has progressed from BM25 (10.1561/1500000019, ) and TF-IDF (aizawa2003information, ) to dense-vector-based (wang2025universal, ; wang2025mf, ; wang2019mcne, ; wang2021hypersorec, ; wang2021decoupled, ) approaches, such as DPR (karpukhin-etal-2020-dense, ) and SEAL (DBLP:conf/nips/BevilacquaOLY0P22, ). Concurrently, reader models have diversified, spanning from extractive readers like DPR-reader (cui-emnlp2019-cmrc2018, ) to generative readers, such as FiD (DBLP:journals/corr/abs-2007-01282, ) and RAG (DBLP:conf/nips/LewisPPPKGKLYR020, ), targeting span and free answering tasks respectively. Recently, the rapid development of Large Language Models (LLMs) (gu2025rapid, ; yin2024entropy, ; shen2024exploring, ; huang2024chemeval, ) has motivated researchers to incorporate them into OpenQA, driven by the models’ advanced abilities in natural language reasoning. For instance, ICL (li2023selfprompting, ) contains to use certain prompts towards better integration of external knowledge with in-context learning, while Replug (shi2023replug, ) enhances an LLM with retrieved knowledge as prompts.
@@
However, when utilizing language models for OpenQA, two notable error phenomena persist, as depicted in Figure 1. These errors include: (1) Knowledge Deficiency. When prompted to identify “which nuclear plant’s detectors first alerted the world to the Chornobyl disaster”, the LLM fabricated an incorrect response. This issue stems from the limited memorization capacity of LLMs. Specifically, LLMs not only face challenges in retaining less frequent knowledge (kandpal2023large, ), but also tend to produce hallucinations (DBLP:conf/emnlp/0001PCKW21, ) when the requisite knowledge is absent from their pretraining data. Existing methods primarily address this issue through pre-training (DBLP:journals/corr/abs-2204-02311, ) or prompting (ren2023investigating, ). Pre-training is too broad and lacks precision while prompting overlooks the model’s potent knowledge storage capabilities (petroni-etal-2019-language, ). (2) Answer Format Alignment. Various OpenQA datasets demonstrate unique answer formats, exemplified by the three markedly different formats depicted in Figure 1. However, even when tasked to produce answers in a specific format, the outputs of LLMs often diverge from these expected formats. This discrepancy is largely due to the distributional biases present in the pre-training corpora, as discussed in (kandpal2023large, ).
@@
Presently, to tackle challenge (1), PaLM 540B (DBLP:journals/corr/abs-2204-02311, ) enlarges the model scale to accommodate a greater wealth of knowledge, which, however, leads to unacceptable computational complexity. So the ideology of RAG (gao2023retrieval, ) integrating knowledge with prompts for efficiency has raised. After researchers recognized the significance of retrieving knowledge, they sought to enhance the quality of retrieved content. Typical examples are enabling large models to autonomously assess the quality of retrieved content for answering questions (ren2023investigating, ), or iteratively conducting multi-step retrievals to progressively enhance the quality of retrieved content and align with the preferences of LLM(asai2023self, ). Nevertheless, integrating knowledge through prompt-based approaches treats LLMs merely as instruction interpreters, disregarding their potent knowledge storage capabilities (petroni-etal-2019-language, ). Consequently, this approach yields unstable and suboptimal results. Toward challenge (2), InstructGPT (ouyang2022training, ) and FLAN (DBLP:conf/iclr/WeiBZGYLDDL22, ) adopt a pre-training and instruction tuning paradigm and attempt to encapsulate all instructions along with the language understanding capabilities equipped with knowledge at the same time. Nevertheless, this paradigm will require the storage of massive real-world data (shen2024optimizing, ; yin2024dataset, ; zhang2025td3, ), and the computational burden is quadratic with respect to the number of model parameters (hoffmann2022training, ). To address this issue, CoF-CoT (nguyen2023cof, ) attempts to utilize a chain of instruction to control the output format. However, a distribution gap persists between the formatted output and knowledge. Consequently, the collaborative training procedure is prone to encountering contradictions (DBLP:conf/acl/MallenAZDKH23, ).
@@
Towards these challenges, in this paper, we propose a novel framework named GenKI, to improve the OpenQA performance by exploring Knowledge Integration and controllable Generation on LLMs. We introduced a novel three-stage paradigm for retrieval, knowledge integration, and controllable generation instead of the two-stage paradigm in RAG, ensuring that the model focuses on either knowledge integration or controlled generation in each instance, to avoid the distributional differences brought by these two tasks. Specifically, in the knowledge integration model, rather than adopting on normal fine-tuning-based method, we propose an innovative method of LLM combining autoregressive training loss and supervised fine-tuning loss to involve the retrieval results. In comparison to the prior research on in-context learning using RAG or fine-tuning methods, which solely instruct the LLM, our model acquires new knowledge more stably by storing the retrieved domain knowledge within parameters rather than relying on prompts. In the controllable generation model, we first utilize a certain fine-tuned LLM to post-process the generated answer towards a certain format. Subsequently, we propose a novel ensemble methodology based on text consistency to incorporate both coherence and fluency from the reward model and faculty from the external selection, thereby ensuring adaptive alignment between the output of our model and the target format.
@@ 2.1. Open-domain Question Answering
OpenQA is a crucial task in NLP and web application (zhang2025killingbirdsstoneunifying, ; ye2025fuxi, ; xie2024breaking, ; xu2024multi, ; xie2024bridging, ; yin2023apgl4sr, ; wang2024denoising, ; han2024efficient, ; han2023guesr, ), which aims to answer information-seeking questions based on auxiliary knowledge source (joshi-etal-2017-triviaqa, ). Traditionally, an OpenQA approach entails a retriever-reader framework, with the retriever used to locate relevant knowledge and the reader generating the response. Extensive research has delved into retriever techniques, encompassing a range from traditional ones such as BM25 (10.1561/1500000019, ) and TF-IDF (aizawa2003information, ) to dense-vector-based methods like DPR (karpukhin-etal-2020-dense, ), as well as retrievers like SEAL (DBLP:conf/nips/BevilacquaOLY0P22, ), which leverages all n-grams in a passage as its identifiers. Since then Iterative retrievers such as GRAPHRETRIEVER (DBLP:journals/corr/abs-1911-03868, ) and GAR (mao-etal-2021-generation, ) have been raised for answering complex questions like those requiring multi-hop reasoning task (yang2018hotpotqa, ). Reader models exhibit a spectrum ranging from extractive readers to generative readers. Extractive readers develop from origin DPR-reader (cui-emnlp2019-cmrc2018, ) to graph-based readers like GRAPHREADER (DBLP:journals/corr/abs-1911-03868, ) and readers ensembling methods like BERTserini (DBLP:conf/naacl/YangXLLTXLL19, ) focus on span-based text answering (cui-emnlp2019-cmrc2018, ). In contrast, generative readers such as FiD (DBLP:journals/corr/abs-2007-01282, ) and RAG (DBLP:conf/nips/LewisPPPKGKLYR020, ) focus on free-form text answering (bajaj2018ms, ; joshi-etal-2017-triviaqa, ). Certainly, given the impressive capabilities showcased by LLMs, developed after generative readers, traditional approaches have witnessed a shift in prominence, often assuming supplementary roles alongside these models. However, valuable methods can still be gleaned from these works. In our study, one approach, namely Dense Passage Retrieval (DPR), was integrated as a retrieval plugin.
@@ 2.2. LLMs-based Open-domain Question Answering
Recent years have witnessed a surge in interest in LLMs for their remarkable capabilities across various NLP tasks like QA and NER (min2021recent, ). Amidst this trend, handling OpenQA based on LLMs is emerging as a popular research direction. Research in this domain can be mainly divided into two categories: discriminative language models-based approaches and generative language models-based approaches. In the first type of approaches, researchers typically fine-tune BERT (devlin2018bert, ) or RoBERTa (liu2019roberta, ) to build a reader aligned to answering tasks (joshi2020spanbert, ; xiao-etal-2021-ernie, ; yasunaga-etal-2022-linkbert, ). Following the occurrence of generative language models, such as GPT (brown2020language, ), GLM (du2021glm, ), LLaMA (touvron2023llama, ), the researchers began adopting these models for OpenQA. This shift was driven by their proven capability to handle OpenQA without relying on retrievers (wei2021finetuned, ). Nevertheless, Self-Prompting (li2023selfprompting, ) persists in trying to enhance LLM itself by in-context learning. There is also some research exploring the role of retrievers in enhancing the performance of generative large models on these tasks, such as REPLUG (shi2023replug, ) and kNN-LM (DBLP:conf/emnlp/ZhongLC22, ). Afterward, researchers recognized the effectiveness of the paradigm using retrieved knowledge, i.e., RAG (gao2023retrieval, ), and began to work on further refining this paradigm.
@@ 2.3. RAG-LLMs on Open-domain Question Answering
Advanced RAG (gao2023retrieval, ) has made improvements to overcome the deficiencies of Naive RAG. The current emphasis of the Advanced RAG model is on enhancing the Retriever and Generator components. On the Retriever enhancement, LangChain (topsakal2023creating, ) first proposed a modularized retrieval and reranking framework. RRR (ma2023query, ) and AAR (yu2023augmentation, ) try to align Retriever with LLM using modularized or end-to-end structure. Finally, the LLaMA-index [^1] https://www.llamaindex.ai proposed an end-to-end alignment and rerank structure to merge them. On the Generator enhancement, end-to-end frameworks (sachan2021end, ) and Pre-training methods like PaLM (chowdhery2022palm, ) are proposed for pre-training LLMs adapting to downstream knowledge. Nevertheless, with the expanding scale of large language models, the viability of the pre-training method diminishes over time. So recently prompt-engineering (li2023selfprompting, ) and supervised fine-tuning(ouyang2022training, ) are the technologies widely used. RFiD (wang2023rfid, ) finetunes the decoder to generate answers by relationships and features. GAG (abdallah2023generator, ) finetunes the model for both answering and generating contextually rich documents tailored to the given question. Self-rag (asai2023self, ) and SeaKR (yao2024seakrselfawareknowledgeretrieval, ) further finetunes an LLM generates and reflects on retrieved passages and its own generations using special retrieval or rerank tokens. However, these generator-based methods merely input retrieved knowledge, which is more precise than pre-training, as prompts into the model. As mentioned in our introduction, the knowledge provided by prompts is not stable. Overall, although the aforementioned generator enhancement methods have achieved remarkable progress in the OpenQA field, current pre-training or fine-tuning-based methods only focus on either full knowledge adaption or applying LLM on both reasoning and answer alignment missions. This leads to limitations in sufficient knowledge integration and controllable generation performance in different task situations, which is also what we focus on.
@@ 3.1. OpenQA Problem Definition
Given a domain knowledge base (e.g., Wikipedia) comprising a set of sentences $P=\{p_{1},p_{2},...,p_{n_{1}}\}$, a specified answer format request $F$, and each query question $q_{i}$ within $Q=\{q_{1},q_{2},...,q_{n_{2}}\}$, OpenQA aims to train a model $M$ that can extract relevant knowledge and generate the ideal answer $a_{i}=M(q_{i},F,P)$ in response to each query, guided by the information in the knowledge base.
In authentic knowledge bases like Wikipedia, information is typically structured within web pages or documents. By segmenting it into sentence form, we can derive the set of sentences denoted as $P$. Take, for instance, the question $q_{i}$: “Where was the initial awareness of the Chernobyl incident triggered?”. Given that the desired format $F$ is Entity Style, the model is expected to leverage insights extracted from these sentences to generate an answer such that $a_{i}=M(q_{i},F,P)=$Sweden”.
@@ 3.2. Instruct Tuning on LLMs
In the OpenQA pipeline with LLMs, we employ a fine-tuned LLM denoted as the model $M$ for answer formulation. The fine-tuning process of the LLM facilitates its adaptation to the distribution and domain knowledge relevant to downstream tasks. This process essentially encompasses an equivalent final objective loss, resembling that of autoregressive training, which is outlined as follows:
(1) $max_{\Phi}\sum_{x_{i},a_{i}\in T}\sum_{t=1}^{|a_{i}|}log(P_{\Phi}(a_{i,t}|x,a_{i,<t})),$
where $\Phi$ represents the parameters of LLM to be optimized, $T$ denotes the training set, $x$ refers to the input context encompassing both an instruction and a query question, and $a_{i,t}$ is the $t$-th token of the generated answer word. For all tunings of our experiment, we adopted the LoRA (hu2021lora, ) method of lightweight fine-tuning by reducing the required GPU memory consumption.

### v2

# Prompting is not Enough: Exploring Knowledge Integration and Controllable Generation on Large Language Models

@@ Abstract.
Open-Domain Quesion Answering, Large Language Model
copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Open-domain question answering (OpenQA) has garnered significant attention in the Natural Language Processing (NLP) community, as it offers enhanced user-friendliness and efficiency compared to conventional search engines. OpenQA seeks to respond to questions utilizing auxiliary knowledge sources (zhu2021retrieving, ), necessitating models that possess both knowledge capability and comprehension abilities. Conventional methods usually entail a retriever-reader framework (DBLP:conf/iclr/DasDZM19, ). The evolution of retriever models has progressed from BM25 (10.1561/1500000019, ) and TF-IDF (aizawa2003information, ) to dense-vector-based approaches, such as DPR (karpukhin-etal-2020-dense, ) and SEAL (DBLP:conf/nips/BevilacquaOLY0P22, ). Concurrently, reader models have diversified, spanning from extractive readers like DPR-reader (cui-emnlp2019-cmrc2018, ) to generative readers, such as FiD (DBLP:journals/corr/abs-2007-01282, ) and RAG (DBLP:conf/nips/LewisPPPKGKLYR020, ), targeting span and free answering tasks respectively. Recently, the rapid development of Large Language Models (LLMs) has motivated researchers to incorporate them into OpenQA, driven by the models’ advanced abilities in recommendation (wang2021hypersorec, ) and natural language reasoning (shen2024exploring, ). For instance, ICL (li2023selfprompting, ) uses certain prompts towards better integration of external knowledge with in-context learning, while Replug (shi2023replug, ) enhances an LLM with retrieved knowledge as prompts.
@@
However, when utilizing language models for OpenQA, two notable error phenomena persist, as depicted in Fig. 1. These challenges include: Challenge (1): knowledge deficiency. When prompted to identify “which nuclear plant’s detectors first alerted the world to the Chornobyl disaster”, the LLM fabricates an incorrect response. This issue stems from the limited memorization capacity of LLMs. Specifically, LLMs not only face challenges in retaining less frequent knowledge (kandpal2023large, ), but also tend to produce hallucinations (DBLP:conf/emnlp/0001PCKW21, ) when the requisite knowledge is absent from their pretraining data. Existing methods primarily address this issue through pre-training (DBLP:journals/corr/abs-2204-02311, ) or prompting (ren2023investigating, ). Pre-training is too broad and lacks precision while prompting overlooks the model’s potent knowledge storage capabilities (petroni-etal-2019-language, ). Challenge (2): answer format alignment. Various OpenQA datasets demonstrate unique answer formats, exemplified by the three markedly different formats depicted in Fig. 1. However, even when tasked to produce answers in a specific format, the outputs of LLMs often diverge from these expected formats. This discrepancy is largely due to the distributional biases present in the pre-training corpora, as discussed in (kandpal2023large, ).
@@
Presently, to tackle challenge (1), PaLM 540B (DBLP:journals/corr/abs-2204-02311, ) enlarges the model scale to accommodate a greater wealth of knowledge, which, however, leads to unacceptable computational complexity. So the ideology of RAG (gao2023retrieval, ) integrating knowledge with prompts for efficiency has raised. After researchers recognize the significance of retrieving knowledge, they seek to enhance the quality of retrieved content. Typical examples are enabling large models to autonomously assess the quality of retrieved content for answering questions (ren2023investigating, ), or iteratively conducting multi-step retrievals to progressively enhance the quality of retrieved content and align with the preferences of LLM(asai2023self, ). Nevertheless, integrating knowledge through prompt-based approaches treats LLMs merely as instruction interpreters, disregarding their potent knowledge storage capabilities (petroni-etal-2019-language, ). Consequently, this approach yields unstable and suboptimal results. Toward challenge (2), InstructGPT (ouyang2022training, ) and FLAN (DBLP:conf/iclr/WeiBZGYLDDL22, ) adopt a pre-training and instruction tuning paradigm and attempt to encapsulate all instructions along with the language understanding capabilities equipped with knowledge at the same time. Nevertheless, this paradigm requires the storage of massive real-world data (yin2024dataset, ), and the computational burden is quadratic with respect to the number of model parameters (hoffmann2022training, ). To address this issue, CoF-CoT (nguyen2023cof, ) attempts to utilize a chain of instruction to control the output format. However, a distribution gap persists between the formatted output and knowledge. Consequently, the collaborative training procedure is prone to encountering contradictions (DBLP:conf/acl/MallenAZDKH23, ).
@@
Towards these challenges, in this paper, we propose a novel framework to improve the OpenQA performance by exploring knowledge integration and controllable generation on LLMs, namely GenKI. We introduced a novel three-stage paradigm for retrieval, knowledge integration, and controllable generation instead of the two-stage paradigm in Retrieval and Generation (RAG), ensuring that the model focuses on either knowledge integration or controlled generation in each instance, to avoid the distributional differences brought by these two tasks. Specifically, in the knowledge integration model, rather than adopting normal fine-tuning-based method, we propose an innovative method of LLM combining autoregressive training loss and supervised fine-tuning loss to involve the retrieval results. In comparison to the prior research on in-context learning using RAG or fine-tuning methods, which solely instruct the LLM, our model acquires new knowledge more stably by storing the retrieved domain knowledge within parameters rather than relying on prompts. In the controllable generation model, we first utilize a certain fine-tuned LLM to post-process the generated answer towards a certain format. Subsequently, we propose a novel ensemble methodology based on text consistency to incorporate both coherence and fluency from the reward model and faculty from the external selection, thereby ensuring adaptive alignment between the output of our model and the target format.
@@ 2.1. Open-domain Question Answering
OpenQA is a crucial task in NLP, which aims to answer information-seeking questions based on auxiliary knowledge source (gu2025rapid, ). Traditionally, an OpenQA approach entails a retriever-reader framework, with the retriever used to locate relevant knowledge and the reader generating the response. Extensive research has delved into retriever techniques, encompassing a range from traditional ones such as BM25 (10.1561/1500000019, ) and TF-IDF (aizawa2003information, ) to dense-vector-based methods like DPR (karpukhin-etal-2020-dense, ), as well as retrievers like SEAL (DBLP:conf/nips/BevilacquaOLY0P22, ), which leverages all n-grams in a passage as its identifiers. Since then Iterative retrievers such as GRAPHRETRIEVER (DBLP:journals/corr/abs-1911-03868, ) and GAR (mao-etal-2021-generation, ) have been raised for answering complex questions like those requiring multi-hop reasoning task (yang2018hotpotqa, ). Reader models exhibit a spectrum ranging from extractive readers to generative readers. Extractive readers develop from origin DPR-reader (cui-emnlp2019-cmrc2018, ) to graph-based readers like GRAPHREADER (DBLP:journals/corr/abs-1911-03868, ) and readers ensembling methods like BERTserini (DBLP:conf/naacl/YangXLLTXLL19, ) focus on span-based text answering (cui-emnlp2019-cmrc2018, ). In contrast, generative readers such as FiD (DBLP:journals/corr/abs-2007-01282, ) and RAG (DBLP:conf/nips/LewisPPPKGKLYR020, ) focus on free-form text answering (bajaj2018ms, ; joshi-etal-2017-triviaqa, ). Certainly, given the impressive capabilities showcased by LLMs, developed after generative readers, traditional approaches have witnessed a shift in prominence, often assuming supplementary roles alongside these models. However, valuable methods can still be gleaned from these works. In our study, one approach, namely Dense Passage Retrieval (DPR), was integrated as a retrieval plugin.
@@ 2.2. LLMs-based Open-domain Question Answering
Recent years have witnessed a surge in interest in LLMs for their scaling (shen2025optimizingsequentialrecommendationmodels) capabilities across various NLP tasks like Recommendation (wu2024survey, ) and NER (min2021recent, ). Amidst this trend, handling OpenQA based on LLMs is emerging as a popular research direction. Research in this domain can be mainly divided into two categories: discriminative language models-based approaches and generative language models-based approaches. In the first type of approaches, researchers typically fine-tune BERT (devlin2018bert, ) or RoBERTa (liu2019roberta, ) to build a reader aligned to answering tasks (joshi2020spanbert, ; xiao-etal-2021-ernie, ; yasunaga-etal-2022-linkbert, ). Following the occurrence of generative language models, such as GPT (brown2020language, ), GLM (du2021glm, ), LLaMA (touvron2023llama, ), the researchers began adopting these models for OpenQA. This shift was driven by their proven capability to handle OpenQA without relying on retrievers (wei2021finetuned, ). Nevertheless, Self-Prompting (li2023selfprompting, ) persists in trying to enhance LLM itself by in-context learning. There is also some research exploring the role of retrievers in enhancing the performance of generative large models on these tasks, such as REPLUG (shi2023replug, ) and kNN-LM (DBLP:conf/emnlp/ZhongLC22, ). Afterward, researchers recognized the effectiveness of the paradigm using retrieved knowledge, i.e., RAG (gao2023retrieval, ), and began to work on further refining this paradigm.
@@ 2.3. RAG-LLMs on Open-domain Question Answering
Advanced RAG (gao2023retrieval, ; jiang2024enhancing) has made improvements to overcome the deficiencies of Naive RAG. The current emphasis of the advanced RAG model is on enhancing the retriever and generator components. On the retriever enhancement, LangChain (topsakal2023creating, ) first proposes a modularized retrieval and reranking framework. RRR (ma2023query, ) and AAR (yu2023augmentation, ) try to align Retriever with LLM using modularized or end-to-end structure. Finally, the LLaMA-index [^1] https://www.llamaindex.ai proposed an end-to-end alignment and rerank structure to merge them. On the Generator enhancement, end-to-end frameworks (sachan2021end, ) and Pre-training methods like PaLM (chowdhery2022palm, ) are proposed for pre-training LLMs adapting to downstream knowledge. Nevertheless, with the expanding scale of LLMs, the viability of the pre-training method diminishes over time. So recently prompt-engineering (li2023selfprompting, ) and supervised fine-tuning(ouyang2022training, ) are the technologies widely used. RFiD (wang2023rfid, ) finetunes the decoder to generate answers by relationships and features. GAG (abdallah2023generator, ) finetunes the model for both answering and generating contextually rich documents tailored to the given question. Self-rag (asai2023self, ), iGFT (tong2025missteps) and SeaKR (yao2024seakrselfawareknowledgeretrieval, ) further finetune an LLM to generate and reflect on retrieved passages and its own generations using special retrieval or rerank tokens. However, these generator-based methods merely input retrieved knowledge, which is more precise than pre-training, as prompts into the model. As mentioned in our introduction, the knowledge provided by prompts is not stable. Overall, although the aforementioned generator enhancement methods have achieved remarkable progress in the OpenQA field, current pre-training or fine-tuning-based methods only focus on either full knowledge adaption or applying LLM on both reasoning and answer alignment missions. This leads to limitations in sufficient knowledge integration and controllable generation performance in different task situations, which is the area we aim to address.
@@ 3.1. OpenQA Problem Definition
Given a domain knowledge base (e.g., Wikipedia) comprising a set of passagess $P=\{p_{1},p_{2},\ldots,p_{n_{1}}\}$, a specified answer format request $F$, and each query question $q$ within query set $Q=\{q_{1},q_{2},\ldots,q_{n_{2}}\}$ with $n_{1}$ and $n_{2}$ representing numbers of passages $P$ and query set $Q$. OpenQA aims to train a model $M$ that can extract relevant knowledge and generate the ideal answer $a_{i}=M(q_{i},F,P)$ in response to each query $q$ with index $i\in[1,2,\ldots,n_{2}]$ with format $F$, guided by the information in the knowledge base.
In authentic knowledge bases like Wikipedia, information is typically structured within web pages or documents. By segmenting it into sentence form, we can derive the set of sentences denoted as $P$. Take, for instance, the question $q_{i}$ is “Where was the initial awareness of the Chernobyl incident triggered?”. Given that the desired format $F$ is entity style, the model is expected to leverage insights extracted from these sentences to generate an answer such that $a_{i}=M(q_{i},F,P)=$“Sweden”.
@@ 3.2. Instruct tuning on LLMs
In the OpenQA pipeline with LLMs, we employ a fine-tuned LLM denoted as the model $M$ for answer formulation. The fine-tuning process of the LLM facilitates its adaptation to the distribution and domain knowledge relevant to downstream tasks. This process essentially encompasses an equivalent final objective loss, resembling that of autoregressive training, the original training loss $\mathcal{L}_{f}$ is outlined as follows:
(1) $\mathcal{L}_{f}=\mathop{\mathrm{max}}_{\Phi}\sum_{x_{i},a_{i}\in T}\sum_{t=1}^{|a_{i}|}\mathrm{log}(\ \mathop{\mathrm{Prob\ }}_{\Phi}(a_{i}[t]|x,a_{i}[<t])),$
where $\Phi$ represents the parameters of LLM to be optimized, $T$ denotes the training set, $x$ refers to the input context encompassing both an instruction and a query question, $Prob(\cdot)$ denotes the generation probability, $a_{i}[t]$ is the $t$-th token ,$||$ denotes the total number of a set while $a_{i}[<t]$ denotes token sequence before the $t$-th token of the generated answer word. For all tunings of our experiment, we adopt the LoRA (hu2021lora, ) method of lightweight fine-tuning to reduce the required GPU memory consumption.